A massless spring of length l and spring constant k is placed vertically on a table. A ball of mass m is just kept on top of the spring. What is the maximum velocity of the ball?

Grade:12

1 Answers

Vishnutheep B
5 years ago
1) 1/2 kx-------- energy gained by the spriNg.. And this gain in energy is completely used in motion of the ball... i.e.. This energy is completely gets transformed into kinetic energy of the block. 
BY CONSERVATION OF ENERGY, 
1/2Kx2= 1/2mv2
Now just find here, 
V=(K/m)1/2x is the velocity of the ball.. .Hope this helps!! 👍